# Daniel Calegari and Andrea Delgado

### v1

@@ Abstract
[Abstract]Educational Process Mining (EPM) is a data analysis technique that is used to improve educational processes. It is based on Process Mining (PM), which involves gathering records (logs) of events to discover process models and analyze the data from a process-centric perspective. One specific application of EPM is curriculum mining, which focuses on understanding the learning program students follow to achieve educational goals. This is important for institutional curriculum decision-making and quality improvement. Therefore, academic institutions can benefit from organizing the existing techniques, capabilities, and limitations. We conducted a systematic literature review to identify works on applying PM to curricular analysis and provide insights for further research. From the analysis of 22 primary studies, we found that results can be classified into five categories concerning the objectives they pursue: the discovery of educational trajectories, the identification of deviations in the observed behavior of students, the analysis of bottlenecks, the analysis of stopout and dropout problems, and the generation of recommendation. Moreover, we identified some open challenges and opportunities, such as standardizing for replicating studies to perform cross-university curricular analysis and strengthening the connection between PM and data mining for improving curricular analysis.
@@ 1 Introduction
Educational Data Mining (EDM) 15, 33 is a field that deals with the analysis of data from educational information systems such as Learning Management Systems (LMS) and administrative systems. Its primary objective is to answer critical academic questions and enhance educational practices. To achieve this, EDM employs different data mining techniques to gain insights into the educational context 33, e.g., grouping similar students based on their learning and interaction patterns and detecting students with difficulties or irregular learning processes, and also to intervene in such context, e.g., predicting an outcome and recommend a more adequate strategy.
This data-centric view of educational data is refined by the process-centric view provided by Educational Process Mining (EPM) 12. In general terms, Process Mining (PM) 42 allows analyzing the records (logs) of events associated with the execution of educational processes, as depicted in Figure 1. The analysis involves three types of activities: discovery, which uses event logs to create a process model from the causal dependencies between events; conformance, which implies verifying the correspondence of the enacted business processes concerning the expected one (through a reference model); and enhancement, which uses additional information to improve the one provided by the process model, obtaining measures such as the duration of the processes, bottlenecks, or the underuse of resources, among others.
Event logs are extracted from educational information systems and are traditionally organized according to the following principles 42:
  • A process comprises cases (traces) representing a process instance execution from start to finish. Each case has a case ID, allowing one to identify a case among others uniquely. As an example, depicted in Figure 1, it could be possible to consider as a case the approval history of a student, with the student ID identifying each case.
@@
Figure 1: EPM schema (from 7) and structure of an event log (both adapted from 42)
@@
Depending on the definition of a case and its events, there are different perspectives from which the data can be analyzed. As synthesized in previous systematic literature reviews 7, 16, 40, EPM pursues different objectives, e.g., to gain a better understanding of the underlying educational process, to detect collaborative or self-regulated student’s learning behavior and difficulties, and to generate recommendations or advice for students and teachers, among others. Many of these objectives relate to analyzing learner’s behavior in the context of a course, e.g., gathering data from Learning Management Systems (LMS) usage based on actual learner’s activity recorded in the event logs 45. However, stakeholders such as study program coordinators and advisors must consider the institutional viewpoint for curriculum understanding and improvement.
@@
An academic curriculum defines a specific learning program 29. It constrains how students are expected to take the courses 47, e.g., defining that a particular course is compulsory or must be taken before or after another. Depending on the flexibility of the curriculum, students have more or less freedom to attend courses and take exams in different order. Curricular improvements must consider students’ behaviors while trying to achieve target educational goals, such as a low dropout rate and a high percentage of students graduating on time 31.
@@
Curriculum mining/analytics 18 focuses on collecting, analyzing, and visualizing the program- and course-level data, such as program structure and course grading, to support curriculum decision-making and quality improvement specifically. It allows analyzing the strengths and weaknesses of a curriculum by identifying its components, relations, and constraints, evaluating how they fit together, identifying potential and actual problems (e.g., blind spots, bottlenecks, and deviations of students’ trajectories), checking underlying beliefs and assumptions, and determining whether the goals have been met.
Curriculum mining is identified as an application domain of EPM 16. There are general reviews on EDM 15, EPM 7, 16, 40, and PM in LMS 45 that mentioned curricular mining. However, although it is a context of utmost importance for decision-making, no specific literature review exists. It prevents us from having an overview of the existing works (the existing reviews are not comprehensive), the challenges that arise for curricular analysis, and the existing research opportunities.
@@
In this work, we conduct a Systematic Literature Review (SLR) following the methodological guidelines in 21, 22 to identify works on applying PM to curricular analysis. We also classify and analyze the existing work to evaluate the PM scope in this domain and provide insights for further research. In particular, we briefly discuss EDM initiatives that can be connected to improve curricular analysis.
@@
The rest of the paper is organized as follows. In Section 2, we describe the conducted literature review methodology, and in Section 3, we synthesize the findings. Then, in Section 4, we discuss the findings and investigate future research directions. Finally, in Section 6, we provide the conclusions of this work.
@@ 2 Literature review methodology
We conducted a Systematic Literature Review (SLR) based on 21, 22. As depicted in Figure 2, it defined three phases: planning, conducting, and reporting the review, the first two divided into several steps that are summarized next.
@@ 2.1 Planning the review
The need for a systematic review is identified in this phase, and the protocol for conducting the review is defined. The protocol includes the formulation of the research question, the definition of the search terms and string, the search sources to be used, and the inclusion and exclusion criteria that will be applied to identify the primary studies of the review.
@@
What are the research papers on applying PM to curricular analysis (i.e., EPM papers) ?
@@
Search keywords, strings, and sources We defined a search string for the selection of published papers. We search for keywords in existing literature reviews 16, 7 and in some papers referred by them. Since many heterogeneous keywords arise, e.g., curriculum, dropout, bottleneck, registration, and trajectories, we used the general string
@@
Inclusion and exclusion criteria The inclusion criteria defined are: (1) the paper concerns PM to curricular analysis, and (2) it is written in English. The exclusion criteria are as follows: (1) the paper is not electronically available on the web, and (2) the paper refers to some non-peer-reviewed publications, such as technical reports, book chapters, proceedings’ prefaces, and journal editorials. We did not exclude short papers since we were interested in extracting exciting ideas beyond the availability of tools or the detailed implementation of case studies.
@@ 2.2 Conducting the review
Execute the searches The initial search returned 3.251 articles, covering papers until 2th November 2023, as shown in Table 1.
Apply inclusion & exclusion criteria We applied inclusion and exclusion criteria to the obtained papers by reading the title, abstract, and keywords. This refinement gives a primary selection of 80 papers (with duplicates), as shown in Table 1.
Select primary studies We selected the primary studies based on reading the full text of the paper. We reinforced the results by backward searching references, searching for authors’ and conferences’ web pages, Google Scholar, and the DBLP database. It allowed finding other papers marked as “Other” in Table 1, giving 34 papers. After removing duplicates, we finally identified 22 primary studies. In the last column of Table 1, we keep one of the sources for duplicates (Scopus and Web of Science are the sources that have duplicates already in the other sources). Primary studies add the specific curricula analysis perspective to the three existing literature reviews that partially refer to the topic 7, 16, 40.
Extract data We extracted the primary data of each work summarized in Table 2 (ordered by year/author). We also extracted data that allowed us to categorize the works and answer the research questions that will be analyzed in the following sections [^1] List of relevant studies, primary selection, and extracted attributes are provided as supplementary material..
@@
IEEE Xplore 41 17 5 5
ACM Digital Library 339 1 0 0
SpringerLink 2331 10 3 3
ScienceDirect 14 2 0 0
Wiley 147 7 0 0
Scopus 242 33 11 7
@@
19 2023 Hobeck R. Process Mining on Curriculum-Based Study Data: A Case Study at a German University Springer
@@ 3 Data synthesis & results
As can be seen, the topic has been studied regularly for the last ten years. The interest is spread in universities in fourteen countries from Asia, Europe, Oceania, and South and North America. Works have been published primarily at conferences. In this case, until 2015, papers were published in more general venues. Since 2015, papers have been published in more specific EDM and education venues, such as the International Conference on Educational Data Mining and the International Conference on Learning Analytics & Knowledge. Moreover, since 2022, papers have been published in a more dedicated venue: the Education meets Process Mining (EduPM) workshop at the International Conference on Process Mining (ICPM). In the case of papers published in journals, even for recent publications from 2022, it can be seen that there are published in non-specialized journals in the EPM, EDM, or the education domains, but on journals focused on more general domains, such as the Database and Expert Systems Applications, Applied Sciences, and the Turkish Journal of Computer and Mathematics Education.
@@
The main objective addressed in these papers is educational trajectories discovery. It means discovering a process model that reflects the trajectory of students, considering a trajectory as the set of courses, exams, or curriculum milestones taken or approved by students on a given period. It also involves identifying the many different variants of students’ behavior and the students’ characteristics that describe such variants. The discovery of educational trajectories is strongly related to the formal modeling of an academic curriculum from predefined requirements. It considers the curriculum’s structure, prerequisites, and business rules, e.g., courses that cannot be taken in parallel, deadlines for approval, and a maximum number of courses that can be taken in one semester.
@@
Even without deviations from the expected curricular trajectory, there could be problems with students’ progress. Bottleneck analysis focuses on identifying and analyzing causes delaying progress (bottlenecks) on the student’s educational trajectories. The stopout/dropout analysis is a more specific analysis, not focused on delays, but on identifying and analyzing causes that lead to stopout and dropout of students. In both cases, the cause analysis is focused on identifying the courses/exams, students’ characteristics, and behavioral patterns that lead to such problems.
Finally, trajectory/curriculum recommendation is devoted to making personalized recommendations to the student’s educational trajectories or the curricular design to improve the expected results, e.g., to earn a degree. These recommendations are based on the information from the analysis described before.
Table 3 summarizes a categorization of the objectives followed by the primary studies. Works are present in many categories when their focus is spread.
@@
Educational trajectories discovery: discovering a process model that reflects the educational trajectory of students. 13 41 2 4
@@
47 13 39
9 27
30 19
Curricular conformance checking: checking whether the observed students’ educational trajectories match their expected curricula. 10 41 4 29
11 47 6
13 46
30 19
Bottleneck analysis: identifying and analyzing causes delaying progress (bottlenecks) on the student’s educational trajectories. 5 29 2 9
@@
Stopout/dropout analysis: identifying and analyzing causes that lead to stopout and dropout of students. 6 29 35
36 27
@@
Trajectory/curriculum recommendation: making personalized recommendations to the student’s educational trajectories or the curricular design to improve the expected results. 6 29 10
48 25
46 32
@@
Research objectives can take the form of questions. As analyzed in 50, whether the PM methodology proposes to start with concrete questions to answer or if there is an exploratory analysis addressing general process-related and data-related questions, a project iteratively refines and generates domain-specific questions of interest. Some works 39, 29, 11, 13, 47, 38, 30 describe curricular analysis specific questions of interest, e.g.,
@@
  • What is the real academic curriculum (study program)?
  • Is there a typical (or the best) way to traverse the curriculum?
  • How does the students’ progression in their first year impact the final grade and graduation time?
  • Are there patterns that capture significant differences in the behavior of students who graduated vs. those who did not?
  • Are there patterns that capture significant changes in students’ behavior from one generation to the other?
  • Which are the best practices and worst practices followed by students?
  • Do current prerequisites between courses make sense?
  • Are the particular curriculum constraints obeyed?
  • Are there patterns that describe a bottleneck in the curriculum?
  • How likely is it that a student will finish the studies successfully or drop out?
  • For late dropout students, how can the educational trajectories that consider the group of high-failure rate courses the student must retake be described?
  • What is a student’s expected time to finish the degree?
@@
  • Are students following the study program more likely to graduate on time?
  • What are the more advantageous courses to take next?
  • Should students take some trajectory to maximize their Grade Point Average?
  • Can a student expect a similar outcome next semester, or must adjust its commitment, engagement, and overall workload?
  • Is it better to drop a course to manage expectations rather than to (predictably) fail and perhaps have to leave the degree altogether?
Two observations arise from these questions. First, we can identify student-related and curriculum-related questions focused on different stakeholders. For example, “What is a student’s expected time to finish the degree?” is focused on providing the student with knowledge about how much time they have left in their career. On the other hand, “Are students following the study program more likely to graduate on time?” seems to be a similar question, but it is focused on knowing if the study plan design allows achieving the goal, assuming the students’ specific usual behavior. Second, we can notice that the answer to some questions can be directly connected to some of the objectives, e.g., curricular conformance checking is directly associated with answering if the particular curriculum constraints are obeyed. However, other questions enclose a combination of requirements that need to be addressed with a broad perspective, as will be discussed in Section 4.
@@
In general terms, as described in 29, curriculum mining refines the general PM tasks within the educational context: building a curriculum model from the observed student’s behavior (discovery), checking whether this behavior matches the expected curricular definitions (conformance checking), and projecting information extracted from the observed data onto the model to perform further analysis and recommend further actions (extension). Table 4 summarizes the type and perspective each work addresses.
These aspects can be further improved by combining PM with data mining and Artificial Intelligence (AI) techniques as in 46, powering the analysis and providing active feedback to students and program coordinators.

### v2

@@ Abstract
[Abstract]Educational Process Mining (EPM) is a data analysis technique that is used to improve educational processes. It is based on Process Mining (PM), which involves gathering records (logs) of events to discover process models and analyze the data from a process-centric perspective. One specific application of EPM is curriculum mining, which focuses on understanding the learning program students follow to achieve educational goals. This is important for institutional curriculum decision-making and quality improvement. Therefore, academic institutions can benefit from organizing the existing techniques, capabilities, and limitations. We conducted a systematic literature review to identify works on applying PM to curricular analysis and provide insights for further research. We reviewed 27 primary studies published across seven major databases. Our analysis classified these studies into five main research objectives: discovery of educational trajectories, identification of deviations in student behavior, bottleneck analysis, dropout / stopout analysis, and generation of recommendations. Key findings highlight challenges such as standardization to facilitate cross-university analysis, better integration of process and data mining techniques, and improved tools for educational stakeholders. This review provides a comprehensive overview of the current landscape in curricular process mining and outlines specific research opportunities to support more robust and actionable curricular analyses in educational settings.
@@ 1 Introduction
Educational Data Mining (EDM) 17, 39 is a field that deals with the analysis of data from educational information systems such as Learning Management Systems (LMS) and administrative systems. Its primary objective is to answer critical academic questions and improve educational practices. To achieve this, EDM employs different data mining techniques to gain insights into the educational context 39, e.g., grouping similar students based on their learning and interaction patterns and detecting students with difficulties or irregular learning processes, and also to intervene in such context, e.g., predicting an outcome and recommend a more adequate strategy.
This data-centric view of educational data is refined by the process-centric view provided by Educational Process Mining (EPM) 12. In general terms, Process Mining (PM) 48 allows analyzing the records (logs) of events associated with the execution of educational processes, as shown in Figure 1. The analysis involves three types of activities: discovery, which uses event logs to create a process model from the causal dependencies between events; conformance, which implies verifying the correspondence of the enacted processes concerning the expected one (through a reference model); and enhancement, which uses additional information to extend or improve an existing process, e.g., adding new perspectives about the duration of the process, bottlenecks or the underuse of resources, among others.
Event logs are extracted from educational information systems and are traditionally organized according to the following principles 48:
  • A process comprises cases (traces) that represent the execution of a process instance from start to finish. Each case has a case ID, allowing one to identify a case among others uniquely. As an example, shown in Figure 1, it could be possible to consider as a case the approval history of a student, with the student ID identifying each case.
@@
Figure 1: EPM schema and structure of an event log (both adapted from 48)
@@
Depending on the definition of a case and its events, there are different perspectives from which the data can be analyzed. As synthesized in previous systematic literature reviews 7, 18, 46, EPM pursues different objectives, e.g., to gain a better understanding of the underlying educational process, to detect collaborative or self-regulated student’s learning behavior and difficulties, and to generate recommendations or advice for students and teachers, among others. Many of these objectives relate to analyzing learner’s behavior in the context of a course, e.g., gathering data from Learning Management Systems (LMS) usage based on actual learner’s activity recorded in the event logs 52. However, stakeholders such as study program coordinators and advisors must consider the institutional viewpoint for curriculum understanding and improvement.
@@
An academic curriculum defines a specific learning program 32. It constrains how students are expected to take the courses 54, e.g., defining that a particular course is compulsory or must be taken before or after another. Depending on the flexibility of the curriculum, students have more or less freedom to attend courses and take exams in different order. Curricular improvements must consider students’ behaviors while trying to achieve the desired educational goals, such as a low dropout rate and a high percentage of students who graduate on time 34.
@@
Curriculum mining/analytics 20 focuses on collecting, analyzing, and visualizing the program- and course-level data, such as program structure and course grading, to support curriculum decision-making and quality improvement specifically. It allows analyzing the strengths and weaknesses of a curriculum by identifying its components, relations, and constraints, evaluating how they fit together, identifying potential and actual problems (e.g., blind spots, bottlenecks, and deviations of students’ trajectories), checking underlying beliefs and assumptions, and determining whether the goals have been met.
Curriculum mining is identified as an application domain of EPM 18. There are general reviews on EDM 17, EPM 7, 18, 46, and PM in LMS 52 that mentioned curricular mining. While numerous studies have applied EPM techniques to various educational contexts, the specific domain of curricular mining lacks a comprehensive synthesis. This gap hinders the development of standardized methods and limits the scalability of insights across institutions.
@@
In this work, we conduct a Systematic Literature Review (SLR) following the methodological guidelines in 23, 24 to identify works on applying PM to curricular analysis. We also classify and analyze the existing work to evaluate the PM scope in this domain and provide insights for further research. This paper contributes to the field by: (1) synthesizing findings from 27 studies on process mining applications in curricular analysis, (2) identifying five primary objectives pursued in this domain, and (3) proposing a research agenda to address open challenges such as standardization, tool integration, and the alignment of process mining with data mining techniques.
@@
The rest of the paper is organized as follows. In Section 2, we describe the conducted literature review methodology, and in Section 3, we synthesize the findings. Then, in Section 4, we discuss the findings and investigate future research directions. In Section 5, we present the threats to the validity of our review. Finally, in Section 6, we provide the conclusions of this work.
@@ 2 Literature review methodology
We conducted a Systematic Literature Review (SLR) based on 23, 24. As depicted in Figure 2, it defined three phases: planning, conducting, and reporting the review, summarized next.
@@ 2.1 Planning the review
This phase identifies the need for a systematic review and defines the protocol for conducting it. The protocol includes formulating the research question, describing the search terms and string, selecting the search sources to be used, and defining the inclusion and exclusion criteria for identifying the primary studies of the review.
@@
What are the research papers on applying process mining to curricular analysis ?
@@
Search keywords, strings, and sources We defined a search string for the selection of published papers. We search for keywords in existing literature reviews 18, 7 and in some papers referred by them. Since many heterogeneous keywords arise, e.g., curriculum, dropout, bottleneck, registration, and trajectories, we used the general string
@@
Inclusion and exclusion criteria Table 1 shows the inclusion and exclusion criteria we have considered. Since we are interested in extracting all kinds of ideas, even those still incipient, we did not exclude short papers. Moreover, we consider extensions of previous articles to have a complete view of the works.
Table 1: Inclusion and exclusion criteria.
Inclusion Exclusion
(1) The paper concerns curricular analysis. (1) The study is not written in English.
(2) The study employs a PM approach as a central element of its analysis. (2) The paper refers to some non-peer-reviewed publications, such as technical reports, book chapters, proceedings’ prefaces, and editorials.
(3) The paper is not electronically available on the web.
@@ 2.2 Conducting the review
Execute the searches The initial search covered papers until 2th November 2023. We then extended the search covering papers until 7th November 2024, giving a total of 3.481 articles, as shown in Table 2.
Apply inclusion & exclusion criteria We applied inclusion and exclusion criteria (Table 1) by reading the title, abstract, and keywords of each paper. This refinement gives a primary selection of 90 papers (with duplicates), as shown in Table 2.
Select primary studies We selected the primary studies based on reading the full text of the paper. We reinforced the results by backward searching references (snowballing), searching for authors’ and conferences’ web pages, Google Scholar, and the DBLP database. It allowed finding other papers marked as “Other” in Table 2, giving 43 papers. After removing duplicates, we finally identified 27 primary studies. In the last column of Table 2, we keep one of the sources for duplicates (ScienceDirect, Scopus and Web of Science are the sources that have duplicates already in the other sources). The primary studies we selected deal with the specific curricular analysis perspective we want to analyze, which was only partially introduced in three existing literature reviews that partially refer to the topic 7, 18, 46.
Extract data We extracted the primary data of each work summarized in Table 3 (ordered by year/author). We also extracted data that allowed us to categorize the works and answer the research questions that will be analyzed in the following sections.
@@
IEEE Xplore 67 17 5 5
ACM Digital Library 369 3 2 2
SpringerLink 2428 14 7 7
ScienceDirect 20 3 1 0
Wiley 165 7 0 0
Scopus 295 36 14 7
@@
21 2022 Hobeck R. Process Mining on Curriculum-Based Study Data: A Case Study at a German University Springer
@@
15 2024 Diamantini C. Evidence-driven appraisal of students’ careers using process mining: a case study Springer
16 2024 Diamantini C. Understanding the stumbling blocks of Italian higher education system: A process mining approach ACM
35 2024 Puttow L. Towards a labeling method for Education Process Mining and a case study on higher education ACM
37 2024 Rennert C. Evaluation of Study Plans using Partial Orders Other
38 2024 Roepke R. Study path analyses for quality assurance and support of study planning Springer
@@ 3 Data synthesis & results
As can be seen, the topic has been studied regularly for the last ten years. The interest is spread in universities in fourteen countries, including Asia, Europe, Oceania, and South and North America. Works have been published primarily at conferences. In this case, until 2015, papers were published in more general venues. Since 2015, papers have been published in more specific EDM and education venues, such as the International Conference on Educational Data Mining and the International Conference on Learning Analytics & Knowledge. Moreover, since 2022, papers have been published in a more dedicated venue: the Education meets Process Mining (EduPM) workshop at the International Conference on Process Mining (ICPM). In the case of papers published in journals, even for recent publications from 2022, it can be seen that there are published in non-specialized journals in the EPM, EDM, or the education domains, but on journals focused on more general domains, such as the Database and Expert Systems Applications, Applied Sciences, and the Turkish Journal of Computer and Mathematics Education.
@@
The main objective addressed in these papers is educational trajectories discovery. It means discovering a process model that reflects the trajectory of students, considering a trajectory as the set of courses, exams, or curriculum milestones taken or approved by students on a given period. It also involves identifying the many different variants of students’ behavior and the students’ characteristics that describe such variants. The discovery of educational trajectories is strongly related to the formal modeling of an academic curriculum from predefined requirements. This process considers the curriculum’s structure, prerequisites, and business rules, such as courses that cannot be taken in parallel, approval deadlines, and the maximum number of courses that can be taken in one semester.
@@
Even without deviations from the expected curricular trajectory, there could be problems with students’ progress. Bottleneck analysis focuses on identifying and analyzing causes delaying progress (bottlenecks) on the student’s educational trajectories. The stopout/dropout analysis is a more specific analysis, not focused on delays, but on identifying and analyzing causes that lead to stopout and dropout of students. In both cases, predictive analytics can be used to anticipate such problems, and the cause analysis is focused on identifying the courses/exams, students’ characteristics, and behavioral patterns that lead to them.
Finally, trajectory/curriculum recommendation is devoted to making personalized recommendations to the student’s educational trajectories or the curricular design to improve the expected results, e.g., to earn a degree. These recommendations are informed by the insights gathered through trajectory discovery, conformance checking, bottleneck, and dropout analyses, collectively guiding improvements in individual student success and curriculum effectiveness.
@@
Educational trajectories discovery: discovering a process model that reflects the educational trajectory of students. 17 47 2 4
@@
54 13 45
@@
33 21
15 16
@@
Curricular conformance checking: checking whether the observed students’ educational trajectories match their expected curricula. 14 47 4 32
11 54 6
13 53
33 21
15 16
37 38
Bottleneck analysis: identifying and analyzing causes delaying progress (bottlenecks) on the student’s educational trajectories. 5 32 2 9
30 33
Stopout/dropout analysis: identifying and analyzing causes that lead to stopout and dropout of students. 8 32 41
42 30
43 44
15 16
@@
Trajectory/curriculum recommendation: making personalized recommendations to the student’s educational trajectories or the curricular design to improve the expected results. 8 32 10 45
55 28
53 36 38
@@
Research objectives can take the form of questions. As analyzed in 57, whether the PM methodology proposes to start with concrete questions to answer or if there is an exploratory analysis addressing general process-related and data-related questions, a project iteratively refines and generates domain-specific questions of interest. Some works 45, 32, 11, 13, 54, 44, 33, 16 describe curricular analysis specific questions of interest, that can be organized into key themes as depicted in Table 5.
Table 5: Research Questions and Related Process Mining Activities
Question Related Process Mining Activity
Curriculum Structure and Pathways
@@
What is the real academic curriculum (study program)? Discovery – Uncovering the pathways students take in the curriculum, capturing the real academic curriculum as implemented.
Is there a typical or optimal way to traverse the curriculum? Discovery and Enhancement – Identifying common patterns (typical pathways) and suggesting optimal traversal sequences.
Do current prerequisites between courses make sense? Conformance Checking – Verifying if prerequisites match actual patterns of student success and curriculum requirements.
Are specific curriculum constraints being followed? Conformance Checking – Checking student adherence to curriculum constraints, such as prerequisites or course load limits.
Student Behavior Patterns and Progression
Are there significant differences in behavior between students who graduated versus those who did not? Discovery and Enhancement – Comparing trajectories of students to identify critical differences.
Do behavior patterns vary significantly between generations? Discovery and Enhancement – Analyzing cohort-based differences to identify shifts or trends in behavior over time.
Are there identifiable best and worst practices among students regarding course-taking behavior? Discovery and Prescriptive Analytics – Identify patterns and use this to guide students toward best practices.
@@
Are students following the recommended study program more likely to graduate on time? Conformance Checking and Predictive Analytics – Evaluating if adherence to the program correlates with timely graduation.
Impact of Early Progress and Predictive Outcomes
How does students’ progression in their first year impact final grades and time to graduation? Predictive Analytics – Analyzing early progress as a predictor of long-term academic outcomes.
How likely is it that a student will successfully complete their studies or drop out? Predictive Analytics – Estimating graduation or dropout likelihood based on students’ early progress and trajectory patterns.
For students who drop out later, what are the typical trajectories, especially concerning high-failure courses that must be retaken? Discovery and Enhancement – Mapping dropout patterns to identify common sequences and factors, like high-failure courses.
Are there early warning signs or specific progression patterns that indicate whether a student might need additional support? Predictive Analytics – Detecting early signs in student trajectories that predict intervention needs.
Bottlenecks and Performance Optimization
Are there patterns indicating bottlenecks in the curriculum? Discovery and Enhancement – Locating frequent delays or bottlenecks where students struggle within the curriculum.
What are the most advantageous courses for students to take next, based on their progress? Prescriptive Analytics – Recommending next courses for students based on historical patterns and successful pathways.
Are there specific course trajectories that can help students maximize their Grade Point Average (GPA)? Prescriptive Analytics – Advising students on pathways that historically lead to high GPAs.
Time to Completion and Graduation Outcomes
What is a student’s expected time to complete their degree? Predictive Analytics – Estimating completion time based on individual student progress and historical data trends.
How does adherence to the program influence on-time graduation? Conformance Checking and Predictive Analytics – Assessing if following the study program correlates with on-time graduation.
Can students expect similar outcomes the following semester, or should they adjust their workload to improve performance? Predictive Analytics and Prescriptive Analytics – Forecasting next-semester performance and advising on workload adjustments.
Decision-Making Support
Is it better to drop a course rather than predictably fail? Predictive Analytics and Prescriptive Analytics – Using predictions to advise on dropping courses to maintain overall progress.
Should students adjust their workload to improve performance? Predictive Analytics and Prescriptive Analytics – Recommending commitment adjustments based on expected performance.
Two key observations arise from these questions. First, we can distinguish between student-related and curriculum-related questions, each focused on different stakeholders and requiring insights at varying levels of detail. For instance, “What is a student’s expected time to complete their degree?” provides individualized insights for students about their likely time to completion, while “How does adherence to the program influence on-time graduation?” addresses aggregate-level insights aimed at understanding whether the study plan effectively supports timely graduation. This distinction reflects the need for granularity of insights for stakeholders, where some analyses are tailored to individual students while others serve curriculum designers and administrators. Second, while some questions align directly with specific objectives, such as using curricular conformance checking to assess whether curriculum constraints are followed, others require a synthesis of multiple analyses. For instance, identifying best and worst practices in student course-taking behavior would likely involve combining insights from trajectory discovery, bottleneck analysis, and success pattern identification. This need for direct versus derived insights highlights the importance of an integrated analytical approach that combines outputs from multiple techniques.
Additionally, these questions vary in terms of their short-term versus long-term focus, with some questions being short-term and action-oriented (e.g., “Is it better to drop a course rather than predictably fail?”) and others focusing on long-term, structural analyses (e.g., “Are specific curriculum constraints being followed?”). Segmenting questions by timeframe can help prioritize analyses based on urgency or impact. Similarly, predictive questions are aiming to forecast outcomes (e.g., “How likely is it that a student will successfully complete their studies or drop out?”) and descriptive questions explaining past behaviors (e.g., “Are there significant differences in behavior between students who graduated versus those who did not?”). Balancing these descriptive, diagnostic, and predictive insights is essential to understand current and future behaviors comprehensively.
Furthermore, several questions imply a need for intervention versus observation, where some aim to generate actionable recommendations (e.g., “Should students adjust their workload to improve performance?”) and others are purely observational, focusing on understanding current behaviors without prescribing changes (e.g., “Do current prerequisites between courses make sense?”). This distinction helps clarify when insights should lead to specific recommendations.
Another theme is cohort and generation analysis, which is evident in questions that explore behavior changes across different generations (e.g., “Do behavior patterns vary significantly between generations?”). This type of question provides insights into curriculum effectiveness over time, potentially identifying the impacts of policy changes or shifts in student demographics.
Finally, many questions reflect a goal of outcome optimization, focusing on enhancing student outcomes, whether it’s maximizing GPA, reducing time to degree, or minimizing dropout risk. This overarching objective unifies student-centered and curriculum-centered inquiries under the shared purpose of improving educational outcomes.
@@
In general terms, as described in 32, curriculum mining refines the general PM tasks within the educational context: building a curriculum model from the observed student’s behavior (discovery), checking whether this behavior matches the expected curricular definitions (conformance checking), and projecting information extracted from the observed data onto the model to perform further analysis and recommend further actions (extension/enhancement). Beyond traditional PM tasks, curriculum mining often incorporates predictive analytics to anticipate future outcomes (such as predicting dropout risk or delays) and prescriptive analytics to recommend actions (like adjusting course load or selecting specific courses). These elements go beyond static model construction by enabling proactive interventions. Table 5 provides insights into the potential relations between PM tasks and the answer to curriculum mining questions.
Table 6 highlighting various aspects concerning the process mining techniques actually addressed by the studies (if specific information is provided in the papers). It includes the process mining type addressed (i.e., Discovery (D), Conformance (C), Enhancement (E), Prediction/Prescriptive Analytics (P)), and the perspective involved. In terms of event semantics, different studies represent academic processes by individual courses taken by students or milestones, such as the completion of semesters or years, to capture progression patterns. The model types vary among works, including Petri Nets, Directly Follows Graphs (DFG), Dependency Graphs (DG), and other notations. For conformance techniques, alignment-based methods are commonly used, though sequence matching and partial orders are also applied in some cases.
Table 6: Process mining techniques.
Paper PM Type Perspective Event Semantics Model Type Conformance
47 D, C control-flow courses Colored Petri Net Alignment-based
2 D, C, E control-flow, time courses Petri Net Alignment-based
4 D control-flow courses Heuristics Net -
32 D, C, E control-flow, time courses Colored Petri Net Alignment-based
10 D, E control-flow, org. courses Social Network -
11 D, C control-flow courses Heuristic Net LTL-based
6 C control-flow milestones - Sequence matching
54 D, C control-flow courses Colored Petri Net -
13 D, C control-flow courses Petri Net Alignment-based
45 D, P control-flow courses - -
9 D, E control-flow, time courses - -
55 D, P control-flow courses DG -
41 D control-flow courses DFG -
42 D control-flow milestones DFG -
30 D, E control-flow, time courses Petri Net, DG -
43 D control-flow milestones DFG -
44 D control-flow milestones DFG -
21 D, C control-flow milestones DFG, Petri Net, BPMN Alignment-based
28 D, P control-flow courses - -
53 D, C, P control-flow courses DFG, Petri Net, BPMN Alignment-based
33 D, C, P control-flow milestones, exams Petri Net Alignment-based
36 D, P control-flow courses DFG, Decision Tree -
15 D, C, P control-flow courses, milestones Petri Net Cost-based fitness
16 D, C, E control-flow, time courses, milestones Causal Net Alignment-based
35 E control-flow courses DFG -
37 D, C control-flow courses, milestones Workflow Net Partial order alignments
38 D, C, P control-flow courses - -